OpenStudy (anonymous):

Trig / Pre Cal / identities These things are killing me. Please help guide me. Thank you. \[ (\sin^4 x-\cos^4 x) = 2\sin^2 x -1 \] \[ (\sin^2 x+\cos^2 x)(\sin^2 x-\cos^2 x)= 2\sin^2 x -1 \] Am I on the right path?

Nnesha (nnesha):

apply special identity \[\huge\rm sin^2 x + \cos^2 =1 \] solve for sin^2

OpenStudy (phi):

yes that is a good start

OpenStudy (anonymous):

so we have \[ (1)(\sin^2 x-\cos^2 x)= 2\sin^2 x -1 \]

OpenStudy (phi):

I would next replace the cos^2

Nnesha (nnesha):

sin^2 x + cos ^2 = 1 solve this identity for cos^2

OpenStudy (anonymous):

\[ (1)(\sin^2 x + 1-\sin^2 x )= 2\sin^2 x -1 \]

Nnesha (nnesha):

\[ \rm (\sin^2 x-(\color{reD}{{1-sin^2x}}))\] write that in the parentheses and distribute by negative one

OpenStudy (phi):

close, but you lost a minus sign -(1 - sin^2)

OpenStudy (anonymous):

\[ (1)(\sin^2 x - (1-\sin^2 x))= 2\sin^2 x -1 \] Now what?

OpenStudy (phi):

-(a +b) means -1*(a+b) = -1*a + -1*b

OpenStudy (anonymous):

Ah so \[ (1)(\sin^2 x - 1+\sin^2 x )= 2\sin^2 x -1 \] I see it now.

OpenStudy (anonymous):

Thank you both!!!

Nnesha (nnesha):

:-)